OpenStudy (anonymous):

If sue can shovel the driveway in 50 minutes and jim can do it in 60, how long would it take for both of them to do it together?

OpenStudy (anonymous):

you wanna do it the real quick snap way?

OpenStudy (anonymous):

using only the two numbers given to you, and no equation at all compute \[\frac{50\times 60}{50+60}\]

OpenStudy (anonymous):

i can explain why that is the right answer if you like, but that is all you really have to do here

OpenStudy (anonymous):

Would you mind explaining for me please?

OpenStudy (anonymous):

ok sure but first is it clear what we are going to compute at the end of the explanation?

OpenStudy (anonymous):

Sorta...

OpenStudy (anonymous):

\[\frac{50\times 60}{50+60}=\frac{3000}{110}=\frac{300}{11}\] or your can write it as a mixed number \[27\tfrac{3}{11}\]

OpenStudy (anonymous):

now that we see how easy it is, ready for the explanation?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok sue can shovel the driveway in 50 minutes that means her rate is \(\frac{1}{50}\) of the job per minute jim can do it in 60 minutes his rate is \(\frac{1}{60}\) of the job per minute

OpenStudy (anonymous):

their combined rate is what you get when you add them and \[\frac{1}{50}+\frac{1}{60}=\frac{50+60}{50\times 60}\]

OpenStudy (anonymous):

i am not going to compute this number just leave it alone

OpenStudy (anonymous):

then since that is their rate, \(\frac{50+ 60}{50\times 60}\) you want the time it takes to complete one job, in other words you want to solve \[\frac{50+60}{50\times 60}T=1\] for \(T\) and you solve for \(T\) by multiplying by the reciprocal so \[T=\frac{60\times 50}{50+60}\]

OpenStudy (anonymous):

you do more than two of these you can go straight to the answer while all your colleagues are trying to write down some equation

OpenStudy (anonymous):

Wow thank you so much I really appreciate the help!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hope it is more or less clear

OpenStudy (anonymous):

it was very clear,, better than how my teacher explained it

OpenStudy (anonymous):

lol thanks!